OpenStudy (anonymous):

Solve |2y - 2| = y. A. {2/3, 2} B. {2, 3} C. {2, 5} D. {1/3, 2}

OpenStudy (anonymous):

a is correct

OpenStudy (anonymous):

Thanks :)

OpenStudy (nurali):

|2y - 2| = y 2y-2=y or 2y-2=-y 2y-y=2 or 2y+y=2 y=2 0r 3y=2 y=2 or y=2/3